# Further author information: (Send correspondence to Jackson A. Zariski)

### v1

@@ Abstract
The WIYN 3.5m Telescope at Kitt Peak National Observatory hosts a suite of optical and near-infrared instruments, including an extreme precision, optical spectrograph, NEID, built for exoplanet radial velocity studies. In order to achieve sub ms-1 precision, NEID has strict requirements on survey efficiency, stellar image positioning, and guiding performance, which have exceeded the native capabilities of the telescope’s original pointing and tracking system. In order to improve the operational efficiency of the telescope we have developed a novel telescope pointing system, built on a recurrent neural network, that does not rely on the usual pointing models (TPoint or other quasi-physical bases). We discuss the development of this system, how the intrinsic properties of the pointing problem inform our network design, and show preliminary results from our best models. We also discuss plans for the generalization of this framework, so that it can be applied at other sites.
@@
To improve the WIYN’s overall operating efficiency and in particular enable NEID to meet its design targets, we are developing a novel method for generating an adaptive, accurate, pointing model using various neural network architectures. We have built a prototype model based on historical telemetry that can accurately predict telescope positioning within roughly , compared to a typical error closer to with the current system. Moreover, such a machine-learning informed model does not require that any additional observing time be dedicated to data collection or re-calibration, as requisite data is gained through normal operation, and model re-calibration can be done offline during daylight hours on a daily basis.
@@ 3 Acquisition
The WIYN presents a very well-suited environment for the development of a deep learning pointing solution, as we have both positional and environmental data logs dating back to 2014, as well as an already established targeting framework in TPoint to use as a baseline for improvement. The development of this system is built off of the Star Tracker camera, an Allied Vision Technologies GT1920 with a Tele-Xenar 2.2/70 lens fitted with an IR filter and a focal plane of 1936 X 1456 pixels. The field of view of the camera is approximately , with a pixel-size of . The camera’s effective bandpass is 400-700 nm [12]. Mounted at the top-end of the telescope, on a frame supporting the secondary mirror, the camera system uses a World Coordinate System to determine pointing by matching up stars in its field to known objects in astrometric catalogs. However, this ‘real’ pointing location is only derived after the fact, when the telescope mount has already moved into position. Hence, we use the calculated celestial coordinate center from the star track center (in ) as the ground truth positions that we aim to predict with our neural network training. These coordinates are hereafter referred to as the AST (astronomical) coordinates. With the Star Tracker camera as a base, our system will represent the pointing center for the telescope as the coordinate center of the image reported by the camera. Note that ultimately we desire a system that is trained to predict the mount encoder measured horizon coordinates, taking celestial coordinates as input. However, the historical dataset does not include the requisite data to analyze the model performance and errors in this direction, so to validate our method we choose to predict the celestial coordinates.
@@
Though the acquisition data is not a uniform time series (targets are not acquired/reacquired at evenly spaced intervals), over the period of a night we do not expect too much deviation in the engineering and temperature variables we hope to capture in our network. In addition, some of these variables are included as features in our network as well, namely those that are quickly measurable relating to atmospheric conditions. Fig. 4 shows how the inputs to our recurrent network can be unraveled into individual observations.
Coordinate | TCS System | Recurrent Neural Network (NN) |
AST RA (Degrees) | 0.2173 | 0.0704 |
AST Dec (Degrees) | 0.2321 | 0.0416 |
As noted above, because of the nature of the current data available to us, we use the mount encoder horizon coordinates to predict celestial coordinates, in order to analyze the performance of our model relative to the existing telescope pointing solution. Summary metrics for our model’s performance are shown in Table 1. For programming these neural networks, we utilized the TensorFlow package in python which allows for the construction of networks with back-propagation done internally on the back-end[1]. We have validated that our model functions in reverse (i.e. using celestial coordinates to predict mount encoder horizon coordinates), but cannot generate a robust comparison with the existing pointing model in this direction. We illustrate the current performance of our model in Fig. 6. Here, NN is the shorthand nomenclature for our recurrent network, TCS refers to the current pointing system[10], and the ‘astropy’ coordinates are shown for reference as the predictions from the python Astropy coordinate transformation package.[3, 11] In addition, Fig.5 shows how our deep learning framework is less prone to systematic errors in the azimuthal direction. Instead we see more uniform errors, and overall improved accuracy in comparison to the existing pointing solution.
While the errors in our neural network model are still relatively large in absolute terms, for example compared to the maximal precision afforded by the resolution of the star tracker camera (), the historical data logs are missing some relevant data that might lead to an improved solution. For example we do not have data on which instrument was being used for each target, nor which instruments were mounted at the time, which affect the weight and balance of the telescope. Finally, while the Star Tracker camera is aligned to the telescope azimuth and elevation, it is not directly aligned to every instrument’s focal plane (some of which have field derotators), which could introduce additional errors. These features could all be included in future telemetry recorded with an eye towards developing new pointing models, and then easily ingested into the neural network training.

### v2

@@ Abstract
The WIYN 3.5m Telescope at Kitt Peak National Observatory hosts a suite of optical and near-infrared instruments, including an extreme precision, optical spectrograph, NEID, built for exoplanet radial velocity studies. In order to achieve sub ms-1 precision, NEID has strict requirements on survey efficiency, stellar image positioning, and guiding performance, which have exceeded the native capabilities of the telescope’s original pointing and tracking system. In order to improve the operational efficiency of the telescope we have developed a novel telescope pointing system, built on both a recurrent neural network and gradient-boosted tree, that does not rely on the usual pointing models (TPoint or other quasi-physical bases). We discuss the development of this system, how the intrinsic properties of the pointing problem inform our network design, and show preliminary results from our best models. We also discuss plans for the generalization of this framework, so that it can be applied at other sites.
@@
To improve the WIYN’s overall operating efficiency and in particular enable NEID to meet its design targets, we are developing a novel method for generating an adaptive, accurate, pointing model using various neural network architectures. In addition, though part of the broader category of machine-learning methods that are not necessarily ‘deep’, we also employ the use of gradient-boosted trees for the best results. We have built a prototype model based on historical telemetry that can accurately predict of telescope acquisitions within roughly a quarter of an arcminute, compared to a typical error closer to arcseconds with the current system. Moreover, such a machine-learning informed model does not require that any additional observing time be dedicated to data collection or re-calibration, as requisite data is gained through normal operation, and model re-calibration can be done offline during daylight hours on a daily basis.
@@ 2.3 Tuning Hyperparameters
2.4 Gradient-Boosted Trees
In addition to the neural network frameworks described above, we also implement a different type of machine-learning, known as gradient-boosted trees, specifically to aid in the acquisition problem. While technically not deep learning, gradient-boosted trees are still a powerful tool for regression tasks, especially in our case where we seek to be accurate on the order of thousandths of a degree. [1] The process of training these models also shares quite a few similarities to the optimization process undertaken by neural networks. Again, we choose a loss function and seek to optimize various parameters based on the residuals minimizing said loss function. Some of these parameters include how the decision tree splits input data into categories, the size of the weights on final outputs applied in a forest of decision trees, as well as the outputs of the trees themselves. Similar to back propagation, these factors are tuned by examining the residuals left from evaluating the training data and using them to find the gradient of said loss function with respect to the above parameters.
@@ 3 Acquisition
The WIYN presents a very well-suited environment for the development of a deep learning pointing solution, as we have both positional and environmental data logs dating back to 2014, as well as an already established targeting framework in TPoint to use as a baseline for improvement. The development of this system is built off of the Star Tracker camera, an Allied Vision Technologies GT1920 with a Tele-Xenar 2.2/70 lens fitted with an IR filter and a focal plane of 1936 X 1456 pixels. The field of view of the camera is approximately , with a pixel-size of . The camera’s effective bandpass is 400-700 nm [13]. Mounted at the top-end of the telescope, on a frame supporting the secondary mirror, the camera system uses a World Coordinate System to determine pointing by matching up stars in its field to known objects in astrometric catalogs. However, this ‘real’ pointing location is only derived after the fact, when the telescope mount has already moved into position. Hence, we use the calculated celestial coordinate center from the star track center (in ) as the ground truth positions that we aim to predict with our neural network training. These coordinates are hereafter referred to as the XY coordinates. With the Star Tracker camera as a base, our system will represent the pointing center for the telescope as the coordinate center of the image reported by the camera. Note that ultimately we desire a system that is trained to predict the mount encoder measured horizon coordinates, taking celestial coordinates as input. However, the historical dataset does not include the requisite data to analyze the model performance and errors in this direction, so to validate our method we choose to predict the celestial coordinates.
@@
Though the acquisition data is not a uniform time series (targets are not acquired/reacquired at evenly spaced intervals), over the period of a night we do not expect too much deviation in the engineering and temperature variables we hope to capture in our network. In addition, some of these variables are included as features in our network as well, namely those that are quickly measurable relating to atmospheric conditions.
| Coordinate | TCS System | Recurrent Neural Network (NN) | XGBoosted Model |
| XY RA (Arcseconds) | 16.2 | 125.1 | 5.20 |
| XY Dec (Arcseconds) | 13.2 | 97.4 | 1.65 |
As noted above, because of the nature of the current data available to us, we use the mount encoder horizon coordinates to predict celestial coordinates, in order to analyze the performance of our model relative to the existing telescope pointing solution. Summary metrics for our model’s performance are shown in Table 1. For programming these neural networks, we utilized the TensorFlow package in python which allows for the construction of networks with back-propagation done internally on the back-end[2]. For the gradient-boosted tree model, we used the XGBoost package also found in Python[1]. We have validated that our model functions in reverse (i.e. using celestial coordinates to predict mount encoder horizon coordinates), but cannot generate a robust comparison with the existing pointing model in this direction. We illustrate the current performance of our model in Fig. 5. Here, Model is the shorthand nomenclature for best model (XGBoost), TCS refers to the current pointing system[11], and the ‘astropy’ coordinates are shown for reference as the predictions from the python Astropy coordinate transformation package.[4, 12] In addition, Fig.4 shows how our machine-learning framework is less prone to systematic errors in the azimuthal direction. Instead we see more uniform errors, and overall improved accuracy in comparison to the existing pointing solution. While the neural network as of now is not as accurate as our XGBoost model, it is still encouraging that we are able to achieve error on just the order of a few arcminutes using nothing but previously obtained Star Tracker data.
Even more encouraging, our best model has an error level that lies in the maximal precision afforded by the resolution of the star tracker camera (), and the historical data logs are missing some relevant data that might lead to an even more improved solution. For example we do not have data on which instrument was being used for each target, nor which instruments were mounted at the time, which affect the weight and balance of the telescope. Finally, while the Star Tracker camera is aligned to the telescope azimuth and elevation, it is not directly aligned to every instrument’s focal plane (some of which have field derotators), which could introduce additional errors. These features could all be included in future telemetry recorded with an eye towards developing new pointing models, and then easily ingested into the neural network and XGBoost training.